# General relativity and the double-zero eigenvalue

### v1

@@ 2.2 Hyperbolic dynamics
which also depends on some parameters collectively denoted by $\lambda$, as will be the case with most of the gravity problems to be analysed below (see Section 4 for a more mathematically rigorous discussion). In a hyperbolic approach, the system is arranged so that, for some range of $\lambda$, it has a hyperbolic equilibrium at $x_{0}$, i.e., none of the eigenvalues of its linear part $D_{x}f(x_{0},\lambda)$ lie on the imaginary axis (including the origin)333in reality the system may have a number of such equilibria in which case one applies the same method to each one of them.. This equilibrium solution persists as does the nature of its stability for solutions $x$ near $x_{0}$ when moving to nearby systems, which as a result share one of the three possible behaviours in this case, namely,
@@
This simplicity in the resulting behaviour of hyperbolic systems appears because in this case the Jacobian matrix $D_{x}f(x_{0},\lambda_{0})$ is invertible for any $\lambda_{0}$ in the said range. In this case, since $D_{x}f(x_{0},\lambda_{0})\neq 0,$ by the implicit function theorem there is a unique smooth solution in the small for $x(\lambda)$ of the equation $f(x,\lambda)=0$, for $\lambda$ sufficiently close to $\lambda_{0}$, for $x$ close to $x_{0}$. So by continuity of the eigenvalues of $D_{x}f(x,\lambda)$ with respect to $\lambda$ sufficiently close to $\lambda_{0}$, the matrix $D_{x}f(x,\lambda)$ also has no eigenvalues on the imaginary axis.
2.3 Dispersive dynamics
Now suppose one can show that the system (2.1) is structurally unstable (‘nonhyperbolic’, or ‘dispersive’) and has a nonhyperbolic equilibrium at $(x_{0},\lambda_{0})$. Then the non-invertibility of the Jacobian (i.e., $D_{x}f(x_{0},\lambda_{0})=0$) implies a necessary condition for the existence of multiple solutions $x(\lambda)$ of the equation $f(x,\lambda)=0$. Hence, the intriguing possibility arises that there are radically new, essentially nonlinear effects emerging for the system (2.1) near $(x_{0},\lambda_{0})$. This means that new solutions may be created or destroyed, and periodic or quasi-periodic behaviours may arise.
We shall call a point $(x_{0},\lambda_{0})$ with,
@@ 2.3 Dispersive dynamics
a singularity [^4] No relation to ‘spacetime singularities’. We hope that the probably excessive use of this term usually associated with the big bang and black holes in gravity research does not preclude its use also in the present, much more informative, context. It is interesting that in the present context spacetime singularities typically do not arise.. This definition implies that a singularity is just a nonhyperbolic point. We note that a singularity need not be a bifurcation point: If we let $n(\lambda)$ denote the number of solutions $x$ for which $(x,\lambda)$ is a solution of the equation $f(x,\lambda)=0$, we say that $(x_{0},\lambda_{0})$ is a bifurcation point of the system (2.1) if it is a singularity and also $n(\lambda)$ changes as $\lambda$ varies in a neighborhood of $\lambda_{0}$.
@@
Most importantly, the phase portraits of the system upon parameter variation undergo continuous transformations, ‘metamorphoses’ to new forms accompanied with topologically inequivalent types of behaviour. An important lesson to be learnt from this is that for a structurally unstable system of dynamical equations, nearby vector fields can have very different orbit structure. This means that when at a certain parameter value $\lambda_{0}$ the vector field $f(x,\lambda_{0})$ has a nonhyperbolic equilibrium where new solutions are then created as the parameter $\lambda$ is varied, one should not only study the orbit structure of the system $f(x,\lambda_{0})$ near $(x_{0},\lambda_{0})$ as is usually done in gravity problems, but also the local orbit structure of nearby systems, that is the systems $f(x,\lambda)$ for $\lambda$ near $\lambda_{0}$ (this is a very important remark usually ignored, however, cf. [3], p. 224).
@@
For example, consider the simplest case of a bifurcating system where one shows that $D_{x}f(x_{0},\lambda_{0})$ has a single zero eigenvalue and the remaining ones having nonzero real parts. What is then the nature of the equilibrium solution $(x_{0},\lambda_{0})$? In this connection, a notable feature of structurally unstable systems is that the laws governing them may be recast in a ‘pure’ form devoid of any hyperbolic behaviour, where the number of equations is drastically reduced and their dynamics is suitably restricted to be on a phase space of smaller dimensionality, the so-called ‘centre manifold’ (unlike the structurally stable ones where this reduction never occurs). In this simplest case, the centre manifold is 1-dimensional, and the problem for a $k$-dimensional family ($k$ being the components of the parameter $\lambda$) is drastically reduced to evolve on a 1-dimensional centre manifold.

### v2

@@ 2.1.2 Hyperbolic dynamics
which also depends on some parameters collectively denoted by $\lambda$, as will be the case with most of the gravity problems to be analysed below. In a hyperbolic approach, the system is arranged so that, for some range of $\lambda$, it has a hyperbolic equilibrium at $x_{0}$, i.e., none of the eigenvalues of the jacobian of its linear part $D_{x}f(x_{0},\lambda)$ lie on the imaginary axis (including the origin)222in reality the system may have a number of such equilibria in which case one applies the same method to each one of them.. Here and throughout we view the equilibrium as a function of a parameter $\lambda$. As $\lambda$ varies, as we shall see below, any eigenvalue whose real part passes through zero marks a critical value $\lambda_{c}$ at which the equilibrium ceases to be hyperbolic (i.e., becomes ‘marginally’ or ‘neutrally’ stable in other terminology) and typically undergoes a bifurcation. An equilibrium $x_{0}(\lambda)$ is hyperbolic if no eigenvalue $\mu(\lambda)$ of $Df(x_{0}(\lambda))$ lies on the imaginary axis. Equivalently, as $\lambda$ approaches a critical value $\lambda_{c}$ at which the real part $\Re\,\mu(\lambda_{c})=0$, the equilibrium becomes neutrally stable, signalling a loss of hyperbolicity and typically the onset of a bifurcation.
A hyperbolic equilibrium solution persists, as does the nature of its stability for solutions $x$ near $x_{0}$ when moving to nearby systems, which as a result share one of the three possible behaviours in this case, namely,
@@
This simplicity in the resulting behaviour of hyperbolic systems appears because in this case the Jacobian matrix $D_{x}f(x_{0},\lambda_{0})$ is invertible for any $\lambda_{0}$ in the said range. Therefore the implicit function theorem guarantees the existence of an open interval $I\ni\lambda_{0}$ and a unique smooth function
$x\colon I\to\mathbb{R}^{n},\quad x(\lambda_{0})=x_{0},$
such that
$f\bigl{(}x(\lambda),\lambda\bigr{)}=0\quad\text{for all }\lambda\in I.$
So by continuity of the eigenvalues of $D_{x}f(x,\lambda)$ with respect to $\lambda$ sufficiently close to $\lambda_{0}$, the matrix $D_{x}f(x,\lambda)$ also has no eigenvalues on the imaginary axis. As we shall see below, this simplicity does not carry over to the non-hyperbolic case.
2.1.3 Dispersive dynamics
Now suppose one can show that the system (2.1) is structurally unstable (‘nonhyperbolic’, or ‘dispersive’) and has a nonhyperbolic equilibrium at $(x_{0},\lambda_{0})$. Then the non-invertibility of the Jacobian implies a necessary condition for the existence of multiple solutions $x(\lambda)$ of the equation $f(x,\lambda)=0$ (instead of a unique solution as before). To see this, suppose instead that
$\det D_{x}f(x_{0},\lambda_{0})=0,$
so that the Jacobian matrix $D_{x}f(x_{0},\lambda_{0})$ is singular. Then $(x_{0},\lambda_{0})$ is a nonhyperbolic (or dispersive) equilibrium and the implicit function theorem no longer guarantees a unique branch. In particular, the failure of invertibility at $\lambda_{0}$ is a necessary condition for the coexistence of two or more smooth solutions $x(\lambda)$ to the equation $f\bigl{(}x(\lambda),\lambda\bigr{)}=0$ near $\lambda_{0}$, and thus signals a potential local bifurcation.
To provide a solid basis for future intuitive interpretations, we may introduce the ‘landscape’ as the graph of all equilibrium solutions
$\{\,(x,\lambda)\mid f(x,\lambda)=0\},$
in the combined $(x,\lambda)$-space. This set may represent the equilibrium set for a given dynamical system (cf. Section 2.2). For most $\lambda$ this graph looks like a single smooth surface (a unique branch $x(\lambda)$). In this landscape analogy, a non-hyperbolic equilibrium corresponds to a valley floor that has been ‘flattened’ (i.e. the Jacobian becomes singular), and suddenly one sheet of the surface can split into two or more sheets, each sheet is a new equilibrium branch. Instead of a single well with a unique bottom, the flattened region harbors several adjacent basins. A small perturbation can then send the system rolling into any one of these dips – each representing a distinct smooth solution branch $x(\lambda)$. Thus, the singular Jacobian signals the emergence of multiple equilibria.
Hence, the intriguing possibility arises that there are radically new, essentially nonlinear effects emerging for the system (2.1) near $(x_{0},\lambda_{0})$. This means that new solutions may be created or destroyed, and periodic or quasi-periodic behaviours may arise. We shall call a point $(x_{0},\lambda_{0})$ with,
@@ 2.1.3 Dispersive dynamics
a singularity [^3] No relation to ‘spacetime singularities’. We hope that the probably somewhat excessive use of this term usually associated with the big bang and black holes does not preclude its usage also in the present context. It is interesting that in the present context spacetime singularities do not typically arise.. This definition implies that a singularity is just a nonhyperbolic point. We note, however, that a singularity need not be a bifurcation point: If we let $n(\lambda)$ denote the number of solutions $x$ for which $(x,\lambda)$ is a solution of the equation $f(x,\lambda)=0$, we say that $(x_{0},\lambda_{0})$ is a bifurcation point of the system (2.1) if it is a singularity and also $n(\lambda)$ changes as $\lambda$ varies in a neighborhood of $\lambda_{0}$. Intuitively, a bifurcation point is where the landscape not only flattens but actually reshapes: as $\lambda$ passes through $\lambda_{0}$, one valley floor can split into two (or two merge into one), so that the number of available basins – i.e. the number of equilibrium branches – changes abruptly.
@@
Most importantly, in this case the phase portraits of the system upon parameter variation undergo continuous transformations, ‘metamorphoses’ to new forms accompanied by topologically inequivalent types of behaviour. An important lesson to be learnt from this is that for a structurally unstable system of dynamical equations, nearby vector fields can have very different orbit structure. This means that when at a certain parameter value $\lambda_{0}$ the vector field $f(x,\lambda_{0})$ has a nonhyperbolic equilibrium where new solutions are then created as the parameter $\lambda$ is varied, one should not only study the orbit structure of the system $f(x,\lambda_{0})$ near $(x_{0},\lambda_{0})$ as is done in a hyperbolic approach to studying gravity problems, but also the local orbit structure of nearby systems, that is the systems $f(x,\lambda)$ for $\lambda$ near $\lambda_{0}$ (this is a very important remark usually ignored, however, cf. [58], p. 224).
For example, consider the simplest case of a bifurcating system where one shows that $D_{x}f(x_{0},\lambda_{0})$ has a single zero eigenvalue and the remaining ones having nonzero real parts. Intuitively, if we think of eigenvalues as vibration frequencies of the landscape, then a zero eigenvalue is a ‘zero-frequency’ mode, i.e. no oscillation at all. In our terrain picture it corresponds to a perfectly flat direction: no slope to pull the system one way or the other, and no restoring force to push it back. Anything that moves a bit along that direction simply stays where one puts it, reflecting the neutral (centre manifold) dynamics along that flat ridge. (To push this interpretation a bit further (when crossing the imaginary axis), a conjugate pair of purely imaginary eigenvalues indicates neutral, oscillatory dynamics, a center in phase space. In our landscape analogy, this corresponds to a perfectly circular, frictionless bowl: a marble given a small nudge will perpetually orbit around the rim at constant amplitude, neither spiraling in (decaying) nor flying out (growing). In other words, the landscape curvature in orthogonal directions provides just the right ‘centripetal’ pull to sustain continuous, undamped oscillations.)
@@
What is then the nature of the equilibrium solution $(x_{0},\lambda_{0})$? In this connection, a notable feature of structurally unstable systems is that the laws governing them may be recast in a ‘pure’ form devoid of any hyperbolic behaviour. In this case, the number of equations is drastically reduced to only the essential ones, and their dynamics is suitably restricted to be on a phase space of smaller dimensionality, the so-called ‘centre manifold’ (unlike in the structurally stable, hyperbolic class where this reduction never occurs). In this simplest case, the centre manifold is 1-dimensional, and the problem for a $k$-dimensional family ($k$ being the components of the parameter $\lambda$) is drastically reduced to evolve on a 1-dimensional centre manifold. Intuitively, a centre manifold is a ‘flat plateau’ through a non-hyperbolic equilibrium – the collection of directions in which neither strong restoring nor repelling forces act, so trajectories linger there, neither growing nor decaying, so motion is slow and ‘marginal’. By restricting the dynamics to this lower-dimensional plateau, one thus captures the slow, critical modes that actually govern how the system will split into new branches as parameters pass through a bifurcation point.